# Asymptotically Unbiased Synthetic Control Methods by Distribution Matching

### v3

@@ 8.2 Auxiliary Covariates
We can consider a case where auxiliary covariates are observable. For each unit $j\in\mathcal{J}$, let
$X_{j,t}\coloneqq(X_{j,t,1},X_{j,t,1},\dots,X_{j,t,K})\in\mathbb{R}^{K}$
be $K$-dimensional covariates, which is not affected by treatment for all $j\in\mathcal{J}$. Let
^$\widehat{\ell}_{k}(\bm{w})=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left\{X_{0,t,k}-\sum_{j\in\mathcal{J}^{U}}w_{j}X_{j,t,k}\right\}.$
When the auxiliary covariates are observable, we extend the DMSCM estimator as
$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0}}\coloneqq v_{0}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right)\right|+\sum^{K}_{k=1}v_{k}\widehat{\ell}^{2}_{k}(\bm{w}),$
where $\bm{v}\coloneqq(v_{k})_{k\in\{0,1,2,\dots,K\}}$ are some weights such that $\sum^{K}_{k=0}v_{0}=1$.
8.3 Distributional SCMs and Optimal Transports
Gunsilius (2023) establies the framework of distributional SCMs. In the framework, that study assumes the following:
$F^{-1}_{Y_{0,t},N}(q)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F^{-1}_{Y_{j,t},N}(q),\ \ \ \forall q\in(0,1),$
where $F^{-1}_{Y_{j,t},N}(q)$ is the quantile function defined as $F^{-1}_{Y_{j,t},N}(q)\coloneqq\inf_{y\in\mathbb{R}}\big{\{}F_{Y_{j,t},N}(q)\geq q\big{\}}$. Then, Gunsilius (2023) estimates the SC weight to match the distribution under the Wasserstein distance. Their method is called DiSCo. Specifically, they estimate the SC weights as
$\widehat{\bm{w}}^{\mathrm{DiSCo}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\frac{1}{M}\sum^{M}_{m=1}\left|\widehat{F}^{-1}_{Y^{N}_{0,t}}(V_{m})-\sum_{j\in\mathcal{J}^{U}}w_{j}\widehat{F}^{-1}_{Y^{N}_{j,t}}(V_{m})\right|,$
where $\widehat{F}^{-1}_{Y^{N}_{j,t}}(V_{m})$ is an empirical quantile function of $Y^{N}_{j,t}$, and $\{V_{m}\}$ is a set of $M$ independent draws from the uniform distribution with support $[0,1]$.
@@ 8.4 Demeaned Density Matching SCM (D2MSCM)
Next, we consider a case where the target distribution does not follow mixture models, but the deviation is zero when it is marginalized. Instead of (3), We consider the following another mixture model
@@
From the viewpoints of linear factor models 1 and fine-grained models (Shi et al., 2022), this assumption allows us $c_{j}\neq 0$ for some $j\in\mathcal{J}$. Under the previous mixture model, we only consider a case where $c_{j}=0$ for all $j\in\mathcal{J}$, which is restrictive in practice.
@@
To estimate the SC weight, we first estimate $c_{j}$ by taking the average of $(Y_{j,t})_{t\in\mathcal{T}_{0}}$ and subtract it from $Y_{j,t}$. This manipulation corresponds to the demeaned estimation in Ferman and Pinto (2021). Let $h\in(0,1)$ be a parameter determined and fixed by us. Then, we estimate SC weights as
@@
$\widehat{\bm{w}}^{\mathrm{D2MSCM}}_{T_{0}}\coloneqq\left(\widehat{w}^{\mathrm{D2MSCM}}_{j,T_{0}}\right)_{j\in\mathcal{J}^{U}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big{(}Y^{N}_{0,t}-\overline{Y}^{N}_{0}\Big{)}^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big{(}Y^{N}_{j,t}-\overline{Y}^{N}_{0}\Big{)}^{\gamma}\right)\right|.$
We refer to this method as the demeaned DMSCM (D2SCM). By using the D2MSCM estimator $(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0}})_{j\in\mathcal{J}^{U}}$, we estimate the treatment effect as
@@
$\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}\coloneqq Y^{I}_{0,t}-\left(\overline{Y}^{N}_{0}+\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{D2SCM}}_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)\right).$ (8)
9 Experiments
We conduct experiments with simulated and real-world data. Additional results are in Appendix 9.1.
9.1 Simulation Studies
In this section, we perform simulation studies to show how our proposed SCMs compared to traditional SCMs. Let $J\in\{10,30,60\}$ and $G\in\{2,3,5,10,20,30,40,50,60,70,80,90,100\}$, $T_{0}=30$, $T_{1}=100$, and $K=5$. We perform a simulation on SCMs for $p_{0,t}(\bm{z})=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p_{0,t}(\bm{z})$, where $\bm{z}$ is ($K+1$)-dimensional ($6$-dimensional) vector, where the first element corresponds to $Y_{j,t}$ and the rest corresponds to $X_{j,t}$. For each $j\in\mathcal{J}$, let $p_{j,t}$ be a density of a multinomial normal distribution $\mathcal{N}(\bm{\mu}_{j,t},\bm{\sigma}^{2}_{j,t})$, where $\bm{\mu}_{j,t}=(\mu_{j,t,1},\mu_{j,t,2},\mu_{j,t,3},\mu_{j,t,4},\mu_{j,t,5},\mu_{j,t,6})$, and $\bm{\sigma}^{2}_{j,t}$ is a $(6\times 6)$-diagonal matrix with a $(k,k)$-element denoted by $\sigma^{2}_{j,t,k}$. Each mean parameter is drawn from a standard normal; distribution. Each variance parameter is drawn from a uniform distribution with a support $[1,20]$. Let us assume that for each $j\in\mathcal{J}$, $\mu_{j,t,1}=\mu_{j,t,1}+\epsilon_{j,t}$ and $\sigma^{2}_{j,t,1}=\sigma^{2}_{j,t,1}+\widetilde{\epsilon}_{j,t}$, where $\epsilon_{j,t}$ and $\widetilde{\epsilon}_{j,t}$ follow a normal densities $\mathcal{N}(0,10)$, respectively. Let us correct the variance as $\widetilde{\epsilon}_{j,t}=0.1$ if $\widetilde{\epsilon}_{j,t}\leq 0.1$ occurs. For the treated unit, we assume that $Y^{N}_{0,t}\sim p_{0,t}(\bm{z})$ for $t\in[t_{0},T_{1}]$ and $Y^{I}_{0,t}=20+Y^{N}_{0,t}$ for $t\in\mathcal{T}_{1}$; that is $\tau_{0,t}=20$. We draw $\bm{w}^{*}$ from a uniform distribution and normalize the sum to be $1$. We apply DMSCM and Abadie $1,000$ times and compute the error of the treatment effect estimation. We show the result with boxplot in Figure 4. Our proposed method successfully estimate the treatment effect with smaller estimation error. The estimation decreases as $G$ increases.
We conduct simulation studies to illustrate the comparative performance of our proposed SCMs and traditional SCMs within a setting analogous to Section 9.1. The only modifications introduced involve the parameters. Specifically, $J\in\{1,5,10,15,20,\dots,35,40,45,50\}$, $G\in\{2,3,5,10\}$, $T_{0}=30$, $T_{1}=1000$, and $K=5$. All other conditions align with those detailed in Section 9.1.
We apply DMSCM and Abadie $1,000$ times and compute the error of the treatment effect estimation in Figure 7. Our focus lies on the fluctuation in mean squared estimation errors as the number of units, $J$, increases. Despite the method by Abadie (2002) incurring escalating estimation errors with the increase of $J$, estimation errors do not exhibit a similar growth under our proposed methods. Furthermore, our proposed methods exhibit lower estimation errors compared to those under the method of Abadie (2002).
A comparable phenomenon is evident in distributional treatment effect estimation. We compare the true and estimated distributions using the Maximum Mean Discrepancy (MMD, Gretton et al., 2012), a probability distance computed through kernel embedding. The upper figure in Figure 8 displays how MMDs alter as $J$ increases. As illustrated, our proposed methods outperform even in this task. The lower figure in Figure 8 presents results for $G\in\{2,10,50,100\}$. As the results indicate, the predictive capability augments as $G$ (the number of moment conditions) increases.
Our findings might have a relationship with those of Spiess et al. (2023), who investigates high-dimensional SCMs. Future investigations will be devoted to exploring these connections.
@@ 9.1 Simulation Studies
Figure 3: Counterfactual outcomes in empirical analysis.
Figure 4: Counterfactual distributions in empirical anaylsis.
Table 2: Results of conformal inference. Estimated treatment effect $\widehat{\tau}_{0,t}$ and $90\%$ confidence interval $(\underline{\tau}_{0,t},\overline{\tau}_{0,t})$.
Terrorist conflict in Basque country Tobacco control program in California Reunification of Germany
Abadie DMSCM Abadie DMSCM Abadie DMSCM
$\gamma$ $2$ $5$ $10$ $100$ $2$ $5$ $10$ $100$ $2$ $5$ $10$ $100$
$\widehat{\tau}_{0,t}$0.800 -0.845 -1.054 -0.279 -0.138 -26.148 -27.744 -32.306 -35.080 -36.879 -1492.374 -1253.616 -1419.120 -1324.448 -510.765
$\overline{\tau}_{0,t}$ 0.522 0.699 0.495 0.521 0.578 8.698 7.852 8.898 9.544 11.748 2635.583 -1008.822 -591.420 -716.376 2327.944
$\underline{\tau}_{0,t}$2.587 -2.150 -2.605 -2.616 -2.590 -36.798 -29.268 -34.783 -37.308 -35.243 -2390.413 -2738.232 -1989.323 -2018.878 -2625.129
@@ 9.2 Empirical Analysis
We conduct empirical studies discussed in Abadie and Gardeazabal (2003), Abadie et al. (2010), and Abadie et al. (2015).
@@
Terrorist conflict in the Basque Country. Abadie and Gardeazabal (2003) investigates the effect of terrorist conflict in the Basque Country on gross domestic product (GDP). In this example, $J=16$, $K=14$, and $(t_{0},T_{0},T_{1})=(1955,1970,1997)$. Basque Euskadi Ta Askatasuna (Basque Homeland and Liberty; ETA) was founded in 1959 as a Basque separatist organization in Spain. ETA promoted an independent Basque state by carrying out terrorist activities mainly in the Basque Country region of Spain. Abadie and Gardeazabal (2003) utilizes annual Spanish regional-level panel data from 1955 to 1997, with a particular focus on the 1970s when ETA conducted significant terrorist activity. In the data, 15 years of pre-intervention data are included. The set of untreated units comprises $17$ other Spanish regions. Each region’s datum contains the outcome of interest (GDP per capita) at the regional level, with covariates including investment rate (percentage), population density, five sectoral productions, and four human capitals, all averaged over the period from 1955 to 1968.
Tobacco control program in California. Abadie et al. (2010) study the effect of a large tobacco control program adopted in California in $1988$, which we explained in our introduction. In this example, $J=20$, $K=5$, and $(t_{0},T_{0},T_{1})=(1970,1989,2000)$. Abadie et al. (2015) investigates the treatment effect of the reunification of Germany. East Germany was absorbed into West Germany, the result being a reunited Germany in 1990. They utilize annual panel data at the country level from 1960 to 2003. The outcome of interest is GDP per capita at the country level. The pre-intervention period spans $30$ years. The set of untreated units comprises $16$ OECD countries that make up synthetic West Germany. Covariates include inflation rate, industry share, investment rate, schooling, and a measure of trade openness.
@@
Reunification of Germany. Abadie et al. (2015) studies the effect of the reunification of Germany on GDP in 1990. Because the countries were dissimilar from any one country to make a comparison group, they apply SCMs to create a composite comparison group. In this example, $J=8$, $K=31$, and $(t_{0},T_{0},T_{1})=(1960,1990,2003)$. Abadie et al. (2010) investigate the causal effect of Proposition 99, a comprehensive tobacco control program adopted by California. They utilize annual state-level panel data from 1970 to 2000. The program was implemented in January 1989, providing 18 years of pre-intervention data. Four states that have adopted large-scale tobacco control programs (Massachusetts, Arizona, Oregon, and Florida) and states that increased state cigarette taxes by $50$ cents or more between 1989 and 2000 (Alaska, Hawaii, Maryland, Michigan, New Jersey, New York, and Washington) and the District of Columbia are excluded from the untreated units of 38 states. The outcome of interest is annual per capita cigarette consumption at the state level, measured as per capita cigarette sales in packs. Covariates include the average retail price of cigarettes, state income per capita (log), percentage of the population aged 15-24, and beer consumption per capita; all averaged over the period from 1980 to 1988.
@@
Results. We show the result of the estimated treatment effect and $90$% confidence interval in Table 2. We show the results of treatment effect estimation in Figure 4 and distributional treatment effect estimation in Figure 4.
@@
For each dataset, we plot the counterfactual outcomes and distributions in Figures 4 and 4, respectively. We also conduct conformal inference on the sharp null hypothesis $H_{0}:\tau_{0,(T_{0}+1)}=\cdots=\tau_{0,T_{1}}=0$ for the treatment effect. For the results of Abadie, DMSCM, We show the $90\%$ confidence interval of the treatment effect in Table 2. Our proposed method has a good fit in terms of the observed outcomes in pretreatment periods $t\in\mathcal{T}_{0}$. Therefore, we consider that our proposed method also predicts counterfactual outcomes and distributions well. In conformal inference, our method does not reject the null hypothesis as with Abadie, except for the DMSCM in the tobacco control problem in California.

### v4

@@ 8.2 Auxiliary Covariates
We can also extend our method to incorporate auxiliary covariates if they are observable. For each unit $j\in\mathcal{J}$, let
$X_{j,t}\coloneqq(X_{j,t,1},X_{j,t,2},\dots,X_{j,t,K})\in\mathbb{R}^{K}$
be $K$-dimensional covariates that are unaffected by treatment. Define
^$\widehat{\ell}_{k}(\bm{w})=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(X_{0,t,k}-\sum_{j\in\mathcal{J}^{U}}w_{j}X_{j,t,k}\right).$
When these covariates are available, we can extend the DMSCM estimator as
$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0}}\coloneqq v_{0}\sum_{\gamma=1}^{G}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|+\sum_{k=1}^{K}v_{k}\widehat{\ell}^{2}_{k}(\bm{w}),$
where $\bm{v}\coloneqq(v_{k})_{k=0}^{K}$ are weights such that $v_{k}\geq 0$ and $\sum_{k=0}^{K}v_{k}=1$.
8.3 Distributional SCMs and Optimal Transport
Gunsilius (2023) propose distributional SCMs, which assume
$F^{-1}_{Y_{0,t},N}(q)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F^{-1}_{Y_{j,t},N}(q)\quad\forall q\in(0,1),$
where $F^{-1}_{Y_{j,t},N}$ is the quantile function defined by
$F^{-1}_{Y_{j,t},N}(q)\coloneqq\inf\left\{y\in\mathbb{R}\colon F_{Y_{j,t},N}(y)\geq q\right\}.$
To estimate $\bm{w}^{*}$, Gunsilius (2023) minimize the Wasserstein distance, referring to their method as DiSCo. Concretely, the DiSCo estimator is
$\widehat{\bm{w}}^{\mathrm{DiSCo}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\frac{1}{M}\sum_{m=1}^{M}\left|\widehat{F}^{-1}_{Y^{N}_{0,t}}(V_{m})-\sum_{j\in\mathcal{J}^{U}}w_{j}\widehat{F}^{-1}_{Y^{N}_{j,t}}(V_{m})\right|,$
where $\widehat{F}^{-1}_{Y^{N}_{j,t}}(V_{m})$ is the empirical quantile function of $Y^{N}_{j,t}$, and $\{V_{m}\}_{m=1}^{M}$ are i.i.d. draws from the uniform distribution on $[0,1]$.
@@ 8.4 Demeaned Density Matching SCM (D2MSCM)
Finally, we consider the case where the target distribution does not precisely satisfy (3), yet any deviation disappears after demeaning. Concretely, instead of (3), we assume
@@
From the perspective of the linear factor model (1) and fine-grained models (Shi et al., 2022), this assumption permits $c_{j}\neq 0$ for some $j\in\mathcal{J}$. Under the previous mixture model, we implicitly required $c_{j}=0$ for all $j\in\mathcal{J}$, which may be restrictive in practice.
@@
To estimate the SC weights in this setting, we first estimate $c_{j}$ by taking the average of $(Y_{j,t})_{t\in\mathcal{T}_{0}}$ and subtracting it from each $Y_{j,t}$. This demeaned approach is inspired by the proposal of Ferman and Pinto (2021) but is applied here to density matching. Let $h\in(0,1)$ be a fixed parameter. We estimate the weights as
$\widehat{\bm{w}}^{\mathrm{D2MSCM}}_{T_{0}}$ $\coloneqq\left(\widehat{w}^{\mathrm{D2MSCM}}_{j,T_{0}}\right)_{j\in\mathcal{J}^{U}}$
@@
$\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma=1}^{G}\frac{2h^{\gamma+1}}{(\gamma+1)!}\biggl{|}\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}-\overline{Y}^{N}_{0}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)^{\gamma}\right)\biggr{|}.$
We call this procedure demeaned DMSCM (D2SCM). Using these weights, we can estimate the treatment effect as
@@
$\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}\coloneqq Y^{I}_{0,t}-\biggl{(}\overline{Y}^{N}_{0}+\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{D2MSCM}}_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)\biggr{)}.$
@@
9 Experiments
We conduct experiments using both simulated and real-world data. Additional results can be found in Appendix 9.1.
9.1 Simulation Studies
We first perform simulation studies to compare our proposed SCMs with traditional SCMs. Set $J\in\{10,30,60\}$ and $G\in\{2,3,5,10,20,30,40,50,60,70,80,90,100\}$, $T_{0}=30$, $T_{1}=100$, and $K=5$.
We consider SCMs for
$p_{0,t}(\bm{z})=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p_{j,t}(\bm{z}),$
where $\bm{z}$ is a $(K+1)$-dimensional (i.e., $6$-dimensional) vector: the first element corresponds to $Y_{j,t}$ and the remaining elements correspond to $X_{j,t}$. For each $j\in\mathcal{J}$, let $p_{j,t}$ be the density of a multivariate normal distribution
$\mathcal{N}\left(\bm{\mu}_{j,t},\bm{\sigma}^{2}_{j,t}\right),$
where $\bm{\mu}_{j,t}=(\mu_{j,t,1},\dots,\mu_{j,t,6})$ and $\bm{\sigma}^{2}_{j,t}$ is a $(6\times 6)$ diagonal matrix whose $(k,k)$-element is $\sigma^{2}_{j,t,k}$. Each $\mu_{j,t,k}$ is drawn from a standard normal distribution, and each $\sigma^{2}_{j,t,k}$ is drawn from a uniform distribution on $[1,20]$. For each $j\in\mathcal{J}$, we modify $\mu_{j,t,1}$ and $\sigma^{2}_{j,t,1}$ by adding noise:
$\mu_{j,t,1}\leftarrow\mu_{j,t,1}+\epsilon_{j,t},\quad\sigma^{2}_{j,t,1}\leftarrow\sigma^{2}_{j,t,1}+\widetilde{\epsilon}_{j,t},$
where $\epsilon_{j,t}$ and $\widetilde{\epsilon}_{j,t}$ follow $\mathcal{N}(0,10)$. If $\widetilde{\epsilon}_{j,t}$ is less than or equal to $0.1$, we set $\widetilde{\epsilon}_{j,t}=0.1$ to avoid near-zero variance.
For the treated unit, we let $Y^{N}_{0,t}\sim p_{0,t}(\bm{z})$ for $t\in[t_{0},T_{1}]$ and set $Y^{I}_{0,t}=20+Y^{N}_{0,t}$ for $t\in\mathcal{T}_{1}$, implying $\tau_{0,t}=20$. We draw $\bm{w}^{*}$ from a uniform distribution over $[0,1]^{J}$ and then normalize it so that its entries sum to $1$.
We apply DMSCM and the method of Abadie and Gardeazabal (2003) (referred to here as Abadie) $1000$ times and evaluate the estimation error of the treatment effect. We plot the estimation error box plots in Figure 3, indicate that our proposed method achieves smaller estimation errors. Moreover, the estimation error declines as $G$ increases.
Next, we consider another set of simulation studies under conditions similar to those above, but with the following changes: $J\in\{1,5,10,15,20,\dots,35,40,45,50\}$, $G\in\{2,3,5,10\}$, $T_{0}=30$, $T_{1}=1000$, and $K=5$. All other settings remain as in Section 9.1.
We again apply DMSCM and Abadie $1000$ times and compute the error of the treatment effect estimation (Figure 3). We focus on how the mean squared estimation error changes as $J$ grows. While the estimation error under the method of Abadie (2002) grows with $J$, our proposed method does not exhibit such growth. Additionally, our method maintains lower estimation errors than the method of Abadie (2002).
A similar pattern appears in distributional treatment effect estimation. We measure the discrepancy between the true and estimated distributions using the Maximum Mean Discrepancy (MMD, Gretton et al., 2012), a probability metric based on kernel embeddings. In the top panel of Figure 4, we show how MMD varies with $J$. As before, our proposed methods outperform the method of Abadie (2002) in this setting. In the bottom panel of Figure 4, we plot MMD for $G\in\{2,10,50,100\}$. As $G$ grows (i.e., as more moment conditions are imposed), the predictive performance improves.
Our observations may connect with findings by Spiess et al. (2023), who examine high-dimensional SCMs. Exploring these connections will be part of our future research.
Figure 5: Counterfactual outcomes in empirical analysis.
Figure 6: Counterfactual distributions in empirical analysis.
Table 2: Results of treatment effect estimation and conformal inference. Estimated treatment effect $\widehat{\tau}_{0,t}$ and $90\%$ confidence interval $(\underline{\tau}_{0,t},\overline{\tau}_{0,t})$.
Terrorist conflict in Basque country.
Abadie DMSCM
$\gamma$- $2$ $5$ $10$ $100$
$\widehat{\tau}_{0,t}$ $-0.800$ $-0.845$ $-1.054$ $-0.279$ $-0.138$
$\overline{\tau}_{0,t}$ $0.522$ $0.699$ $0.495$ $0.521$ $0.578$
$\underline{\tau}_{0,t}$ $-2.587$ $-2.150$ $-2.605$ $-2.616$ $-2.590$
Tobacco control program in California.
Abadie DMSCM
$\gamma$- $2$ $5$ $10$ $100$
$\widehat{\tau}_{0,t}$ $-26.148$ $-27.744$ $-32.306$ $-35.080$ $-36.879$
$\overline{\tau}_{0,t}$ $8.698$ $7.852$ $8.898$ $9.544$ $11.748$
$\underline{\tau}_{0,t}$ $-36.798$ $-29.268$ $-34.783$ $-37.308$ $-35.243$
Reunification of Germany.
Abadie DMSCM
$\gamma$- $2$ $5$ $10$ $100$
$\widehat{\tau}_{0,t}$ $-1492.374$ $-1253.616$ $-1419.120$ $-1324.448$ $-510.765$
$\overline{\tau}_{0,t}$ $2635.583$ $-1008.822$ $-591.420$ $-716.376$ $2327.944$
$\underline{\tau}_{0,t}$ $-2390.413$ $-2738.232$ $-1989.323$ $-2018.878$ $-2625.129$
@@ 9.2 Empirical Analysis
We now present empirical studies based on Abadie and Gardeazabal (2003), Abadie et al. (2010), and Abadie et al. (2015).
@@
Terrorist conflict in the Basque Country. Abadie and Gardeazabal (2003) investigate the effect of terrorist conflict in the Basque Country on gross domestic product (GDP). In this example, $J=16$, $K=14$, and $(t_{0},T_{0},T_{1})=(1955,1970,1997)$. Basque Euskadi Ta Askatasuna (Basque Homeland and Liberty; ETA) was founded in 1959 as a Basque separatist organization in Spain, seeking an independent Basque state primarily through terrorist activities in the Basque Country region. Abadie and Gardeazabal (2003) use annual Spanish regional-level panel data from 1955 to 1997, with a particular focus on the 1970s when ETA’s terrorist activity was significant. The dataset includes fifteen years of pre-intervention observations. The set of untreated units consists of $17$ other Spanish regions. Each region’s record contains the outcome of interest (GDP per capita) and additional covariates (investment rate, population density, five sectoral output measures, and four measures of human capital), all averaged from 1955 to 1968.
Tobacco control program in California. Abadie et al. (2010) study the effect of a large tobacco control program adopted in California in 1988 (described in the Introduction). Here, $J=20$, $K=5$, and $(t_{0},T_{0},T_{1})=(1970,1989,2000)$.
Reunification of Germany. Abadie et al. (2015) examine the impact of Germany’s reunification in 1990. Because no single country provided a suitable comparison, SCMs were used to create a composite unit. In this example, $J=8$, $K=31$, and $(t_{0},T_{0},T_{1})=(1960,1990,2003)$.
@@
Separately, Abadie et al. (2010) also investigate the causal effect of Proposition 99, a comprehensive tobacco control program adopted by California. They use annual state-level panel data from 1970 to 2000. The program was implemented in January 1989, providing eighteen years of pre-intervention observations. Four states that adopted large-scale tobacco control programs (Massachusetts, Arizona, Oregon, and Florida) and states that raised cigarette taxes by at least \$0.50 between 1989 and 2000 (Alaska, Hawaii, Maryland, Michigan, New Jersey, New York, and Washington) as well as the District of Columbia are excluded from the set of 38 control states. The outcome of interest is annual per capita cigarette sales in packs; covariates include the average retail price of cigarettes, log of state income per capita, the percentage of the population aged 15–24, and per capita beer consumption, all averaged between 1980 and 1988.
@@
Results. Table 2 reports the estimated treatment effects and 90% confidence intervals. Figures 6 and 6 show the estimated treatment effects and the distributional treatment effects, respectively.
@@
In each dataset, we plot the counterfactual outcomes and distributions in Figures 6 and 6. We also perform conformal inference on the sharp null hypothesis $H_{0}:\tau_{0,(T_{0}+1)}=\cdots=\tau_{0,T_{1}}=0$ for the treatment effect. For both Abadie’s method and DMSCM, we present the 90% confidence interval of the treatment effect in Table 2. Our proposed method fits the observed outcomes well in the pre-treatment period ($t\in\mathcal{T}_{0}$) and, accordingly, produces plausible counterfactual outcomes and distributions. In the conformal inference analysis, our method does not reject the null hypothesis in the same manner as Abadie’s estimator, except for the DMSCM in the California tobacco control example.